The Lineweaver-Burke Plot for the enzyme in question 1 is shown below. In this plot, the y intercept (0.019153)   is 1/vmax and the x intercept (-0.0449) is -1/km.

  1. What are the exact km and vmax based on this graph (show your work)?

                Km=

                Vmax=
Answer #1

Y Intercept 1/Vmax = 0.019153

Vmax = 1 / 0.019153 = 52.21

Y Intercept -1/Km = -0.0449

Km = 1/ 0.0449 = 22.27